# The Overcooked Generalisation Challenge

### v2

@@ 5 The Overcooked Generalisation Challenge
Figure 3 provides an overview of the Overcooked Generalisation Challenge. The OGC extends previous work by evaluating the cooperative abilities out-of-distribution. In contrast to existing UED environments, the OGC focuses on the cooperation of multiple agents in a complex, cooperative task across different levels. More specifically, two different agents are tasked with cooking a soup together in the five original layouts of Overcooked-AI (see Figure 1 bottom) but without encountering them or their partner during training. The original five layouts have been designed to test and explore different kinds of cooperation and thus form suitable out-of-distribution test levels. Since the OGC does not limit evaluation to one specific physical environment or partner, it is more closely aligned with real-world human-AI collaboration. To train an agent capable of generalisation, we generate a curriculum of training layouts via procedural content generation. To generate a curriculum of layouts, we use DCD methods. Specifically, methods in which an environment designer interacts with the challenge by designing layouts either from scratch through interacting with OvercookedUED – a novel environment for creating Overcooked levels – by alternating existing layouts through the Overcooked mutator or by letting the OGC generate random layouts. At every step of the curriculum, this designer must account for agents’ cooperation ability when trying to generate layouts that are at the forefront of their abilities. While OGC refers to the challenge as a whole, it comprises several components that enable its integration with DCD algorithms (see Figure 3). At the heart, it features an Overcooked environment capable of running different levels fast and in parallel in which agents learn to cooperate. It features OvercookedUED that provide methods, interfaces and a teacher environment to design novel layouts and an Overcooked Mutator that alters existing layouts.
Overcooked-AI
OGC builds on the Overcooked-AI environment provided by the JaxMARL project (Rutherford et al., 2024b), keeping features consistent with the original implementation. This includes action and observation spaces, i.e. the set of actions is {left, right, up, down, interact, stay} and observations are encoded as a stack of 26 $h\times w$ boolean masks encoding the positions of elements in the environment. In this representation, the first mask encodes the position of the first agent; the second mask is the one of the second agent, etc. We extended their Overcooked implementation to be parallelisable over different layouts – which is impossible in the JAXMarl version – to train agents on many different layouts simultaneously and efficiently. This requires padding all layouts to the same size. I.e., during rollouts, layouts are padded to a maximum size, and all objects in these layouts are one-hot encoded based on their position in equally sized masks. While this facilitates fast parallel rollouts that can be compiled just-in-time, it requires the introduction of a maximum height $h$ and width $w$ that need to be picked as a hyper-parameter before training.
OvercookedUED
OvercookedUED features the interfaces necessary to design new layouts. For algorithms that make use of a teacher agent to create layouts (PAIRED, etc.), OvercookedUED provides a teacher environment (see Figure 3 right). This environment allows a teacher policy to take design steps to parametrise the underspecified MDP. At every timestep $t$ of the generation process, the teacher observes the unfinished layout and picks an action from a space that consists of the total number of cells in the $h\times w$ grid. This cell then becomes filled with the next item to be placed. Objects are placed sequentially and in a deterministic order, starting from walls, agents one and two, goal, onion, pot and bowl positions. An episode in the teacher MDP lasts until all items are placed. When placing a new element on top of an existing one of a different type, the new element is placed randomly on a free cell. Placing two elements of the same type on the same cell results in the second not being placed. This allows designers to place various elements per type, e.g. one to two onions. This is consistent with prior work (Dennis et al., 2020). As in previous work (Jiang et al., 2023), OvercookedUED does not check whether a layout is solvable and leaves the task of designing and/or identifying suitable training layouts to the DCD method. For algorithms that do not specify a teacher, such as PLR, OvercookedUED generates random layouts. These random layouts feature one or two piles of onions, bowls, pots and serving locations, and two agents. Finally, some DCD algorithms, such as ACCEL, require alternating existing layouts by mutating them. OvercookedUED supports layout mutation through five basic operations: (1) converting a random wall to a free space and vice versa, (2) moving goals, (3) pots, (4) plate piles, and (5) onion piles. Our mutator does not move agents. Given a layout, our mutator randomly samples $n$ operations and applies them. All versions allow the number of walls to be configured, and the environment always places a border wall.
@@
Figure 4: The SoftMoE-LSTM agents architecture used in this work. We employ the PerConv tokenisation technique introduced in (Obando-Ceron et al., 2024).
Evaluation
We evaluate agents by their performance on out-of-distribution Overcooked-AI layouts to asses generalisation performance in self-play, ad-hoc teamplay (Stone et al., 2010) and zero-shot coordination. In ad-hoc teamplay, a fictitious co-play (Strouse et al., 2021, FCP) and maximum entropy-based population-based training (Zhao et al., 2023, MEP) population of a total 24 agents each is used to asses cooperation performance with novel partners. Both populations are trained with equal settings and include a low, medium and high-skilled checkpoint of each run extracted at 10%, 50%, and 100% achieved return, respectively. The population entropy coefficient $\alpha$ is $0.01$ for MEP. We use cross-play with a diverse population and zero-shot cooperation as a proxy for assessing the abilities of an agent to coordinate with humans. In zero-shot cooperation, we task agents to coordinate with different random seeds of the same method (Hu et al., 2020). Our benchmark includes the necessary tools to perform this evaluation. In our analysis, we report results using the mean episode reward and mean layout solved rate, similar to previous work (Jiang et al., 2023). A layout is considered solved if an agent pair delivers more than one soup, differentiating goal-directed from random behaviour. We present these metrics in the self- and cross-play settings. Additionally, we investigate what kinds of levels agents perform poorly in and why in a final error analysis.
6 Analysing & Benchmarking
@@
Figure 6: Sample levels generated by the different methods after $15,000$ (Middle) and $30,000$ (End) epochs. Even after considerable training, none of the methods can guarantee the generation of solvable layouts (Middle-row leftmost and rightmost).
We establish baseline performance on the OGC using state-of-the-art architectures, DCD algorithms, and neural network architectures. We aim to set a performance baseline for future works and show what evaluations this benchmark enables. To this end, we first show that generalising to novel layouts in Overcooked is difficult, and then we move on to the additional challenge of zero-shot cooperation.
All baselines are trained using MAPPO, which works well in cooperative settings (Yu et al., 2022). As for DCD algorithms, we compare the performance of DR, PLR⟂,∥, Pop. PAIRED and ACCEL. We chose these methods as they have better theoretical guarantees (PLR vs PLR), better runtime performance (ACCEL and PLR), or because we found them to perform better empirically (Pop. PAIRED vs PAIRED). We excluded POET (Wang et al., 2019) in this analysis as it outputs specialists rather than generalists, which we require (Parker-Holder et al., 2022). Additionally, we excluded MAESTRO as it is based on prioritised fictitious self-play (Heinrich et al., 2015; Vinyals et al., 2019) that is not readily adaptable to the cooperative setting (Strouse et al., 2021). In our experiments, the number of walls to place for a specific layout is initially drawn randomly (PLR, ACCEL, DR) or chosen by the action of the teacher agent (PAIRED) from the interval $[1,15]$ where $15$ is the configured maximum. From there on, walls are placed at random (PLR, ACCEL, DR) or based on the actions of the teacher agent (PAIRED), which is parametrised by its own neural network. All methods place at least one and up to two items per category. Since ACCEL can add walls through editing, we restrict ACCEL from adding more walls than the maximum configured to keep methods comparable. In this case, ACCEL can only remove walls or move items. The same holds for PAIRED, where we pick the number of design steps so that no more than the maximum number of walls can be set. The design space of all methods is thus the same.
As in (Jiang et al., 2023), if not stated otherwise, we train in $32$ parallel environments and stop after $30,000$ outer training loops, amounting to just under $400$ million steps in the environment. Hyperparameters were picked after a grid search over reasonable values, and all parameters are provided in Appendix A.4.
Neural Architectures and Agents
Table 1: Mean episode reward for the different methods averaged over the respective testing layouts. The best result is shown in bold. We report aggregate statistics over three random seeds. We include Oracles trained on the five testing layouts to establish an empirical upper bound.
@@
DR $0.46\pm 0.16$ $5.22\pm 7.19$ $00.00\pm 0$ 0.00
PLR $0.17\pm 0.06$ $0.91\pm 0.71$ $0.12\pm 0.15$
PAIRED $0.64\pm 0.38$ $\mathbf{9.52\pm 1.02}$ $00.00\pm 0$ 0.00
ACCEL $0.40\pm 0.35$ $0.72\pm 0.59$ $0.09\pm 0.12$
Oracle $\mathbf{189.49}\pm 12.96$ $\mathbf{217.02}\pm 39.18$ $\mathbf{155.01}\pm 12.82$
@@
We first compared suitable network architectures for our agents. Our default neural network architecture consists of a convolutional encoder with a recurrent neural network with an LSTM (Hochreiter & Schmidhuber, 1997). We picked it because of its performance in previous work (Yu et al., 2023). In addition, we also explore the use of SoftMoE (Obando-Ceron et al., 2024), which have recently been identified for their potential for enabling scaling and generalisation, and S5 (Smith et al., 2023) layers due to the strong results of structured state-space models (Gu et al., 2022) in meta reinforcement learning (Lu et al., 2023). SoftMoE modules replace the penultimate layer after the feature extractor and S5 layers the LSTM in all experiments. For SoftMoE we use the PerConv tokenisation technique (Obando-Ceron et al., 2024). We hypothesise that these provide better generalisation performance. All details are in Appendix A.5.
We first compare their ability to fit all evaluation layouts simultaneously. We refer to these agents as Oracles and use them later for comparison. As presented in Figure 5 (left), all agents can fit all five layouts simultaneously. Note that the S5 agent underperforms considerably while the SoftMoE-LSTM agent performs best, which is why we will use them in future experiments. As a baseline, we compare against a SoftMoE-LSTM agent that is only fitted to one layout and evaluated on all (Figure 5 (right)) which overfits. Because of our finetuned SoftMoE model and implementation, we see very high self-play returns compared to the literature. The SoftMoE-LSTM Oracle sometimes even outperform the overfitting baseline. Detailed results are in Figure 5 and training curves are in Appendix A.7.1.
Layout Generalisation Performance
Despite our fine-tuned implementation, simply generalising to the testing layouts in the OGC is already challenging for all methods without coordinating with novel partners, as presented in Table 1. Compared to commonly used single-agent Maze environments such as AMaze, all DCD methods struggle to obtain good results. This is most evident when compared with oracle policies (bottom row). Since the environment features many moving parts to be placed correctly (see Figure 2), approaches like DR have difficulties in finding optimal placements by chance, as reflected in the results. PAIRED outperforms all other models significantly $p<0.05$ using a one-sided paired t-test. This is also shown in the mean solved rate where it reaches $14.6\pm 7.7\%$, while all other models have a solved rate of mostly $0\%$ (cf. Appendix A.6.2). While this model performs better on average, layouts differ greatly in their difficulty. Our best-performing model reaches modest performance in Asymmetric Advantages and Cramped Room while mostly failing in others, with no other model achieving noteworthy results. We show the training and evaluation return of SoftMoE-LSTM-PAIRED in Figure 7 and 15. The curve shows the model converging but never generalising to the test levels. The full results are in the Appendix A.6.3.
Performance with Novel Partners
@@
Figure 8: Ad-hoc team-play performance of SoftMoE-LSTM-PAIRED and several baselines with both an MEP and an FCP expert population. While SoftMoE-LSTM-PAIRED outperforms simple baselines, it fails to compete against the oracle.
@@
Figure 9: Zero-shot cooperation performance of both the oracle and the SoftMoE-LSTM-PAIRED agents (trained with different seeds).
Ultimately, we want the OGC to connect map and partner generalisation and thus evaluate both ad-hoc team play and zero-shot coordination. For ad-hoc team play, we train and then use an FCP and MEP population to establish how our and future agents can coordinate with diverse policies. See Appendix A.6.4 for details. We present results for the SoftMoE-LSTM-PAIRED policy in Figure 8 together with three other baselines: stay, which is a partner that never moves, random, which samples random actions and oracle as discussed above. Results for all methods are in the Appendix in Figure 13. As performance on out-of-distribution levels rises, agents become more competent at coordinating. PAIRED always outperforms baselines (cf. Appendix A.6.5). However, even PAIRED policies often perform only slightly better than random baselines, which signifies the challenges of our benchmark. This is also evidenced by the kinds of levels these methods generate (Figure 6), as they tend to pivot towards generating open spaces that ease cooperation but are notably different from evaluation layouts. Overall, cooperation performance is carried mainly by the expert FCP and MEP agents (compare Table 13 and 14), mostly since our agents struggle to perform on the evaluation layouts in the first place.
Additionally, to evaluate our agents’ full spectrum of out-of-distribution abilities, we establish zero-shot cooperation performance with agents themselves over three random seeds. We display results in Figure 9. Two results stand out in the Figure. First, our SoftMoE-LSTM-PAIRED agents perform badly in both self- and cross-play in the three harder evaluation layouts (Coordination Ring, Forced Coordination and Counter Circuit). Second, in the two remaining easier layouts, our agent sometimes performs better with a novel partner than with itself. The same is never true for the oracle baseline. While this does not present definite evidence, it hints at a connection between map generalisation and zero-shot cooperation performance in Overcooked.
Error Analysis

### v3

@@ 4 The Overcooked Generalisation Challenge
Figure 3: The OGC features a large design space in which many different elements have to be placed in relation to each other, creating challenging environments for both environment designers and agent training.
The Overcooked Generalisation Challenge is a new benchmark that allows us to evaluate agents on their ability to cooperate with unknown partners in previously unseen environments. Unlike existing benchmarks, the OGC combines unsupervised environment design with multi-agent coordination, introducing the first open-source UED testbed for cooperative RL covering both environment and partner generalisation. Built on Overcooked-AI Carroll et al. (2019), OGC integrates with DCD algorithms to support procedural training, layout mutation, and zero-shot evaluation across complex coordination tasks.
Overcooked-AI is one of the most popular environments for cooperative RL and ad-hoc teamwork. In the environment, two agents are tasked with cooking and delivering a soup together. Delivering a soup entails collecting and putting onions into a pot, letting it cook, collecting the soup with a plate and delivering it to a serving station, which results in a reward of 20.
Figure 2 shows how OGC interfaces with various DCD methods. It supports both teacher-based approaches (e.g., PAIRED (Dennis et al., 2020)) and edit-based methods (e.g., ACCEL (Parker-Holder et al., 2022)) using mutator functions that transform existing layouts. Figure 3 illustrates the complexity of layout design in Overcooked, where task difficulty depends critically on the spatial configuration of multiple interdependent objects and agents.
4.1 Environment and Layout Generation
OGC extends the JaxMARL implementation of Overcooked-AI (Rutherford et al., 2024b), which defines a discrete action space left, right, up, down, interact, stay and an observation space consisting of 26 binary masks of size $h\times w$, encoding the positions of agents, objects, and obstacles. To support large-scale training, we enable parallel rollouts across multiple layouts, requiring all layouts to be padded to a fixed maximum height $h$ and width $w$. This enables fast training and execution speeds across hundreds or thousands of environments using JAX.
4.2 Curriculum Learning in OGC
OGC exposes two core interfaces for DCD methods: OvercookedUED and the Overcooked Mutator.
OvercookedUED implements a teacher environment where a generator policy sequentially places objects onto a layout grid. At each step $t$, the teacher selects a grid cell and places one object from a fixed sequence (walls, agents, goals, ingredient piles, pots, bowls). If the target cell is already occupied, the object is placed randomly in a free cell of the same type. Placing two elements of the same type in the same location results in the second being ignored, enabling variable object counts per type – consistent with prior UED designs (Dennis et al., 2020). For UED methods that lack a teacher component (e.g., PLR), OvercookedUED also provides a random environment generator that follows the same structure as the teacher but samples object positions uniformly.
The Overcooked Mutator enables layout evolution for edit-based methods. It supports five operations: (1) toggling walls and free spaces, (2–5) moving goal, pot, plate, and onion pile positions. Agent start positions remain fixed. The number of mutations applied can be configured to control curriculum granularity.
All versions leave layout solvability unchecked, following the convention in prior UED work (Dennis et al., 2020; Jiang et al., 2023), and place responsibility for level quality on the DCD method.
4.3 Evaluation Protocol
Figure 4: First, we propose to evaluate agents in self-play on the five original layouts and several layouts that are created from several symmetry classes to evaluate their ability to generalise. We combine a range of transformations shown in the bottom row to generate 28 additional layouts. Second, we evaluate coordination with novel partners in the original five.
We study three evaluation modes: Self-play, zero-shot and ad-hoc coordination.
We propose to test agents in self-play to evaluate how well they generalise to novel levels. Figure 4 illustrates our evaluation suite. We use the five original Overcooked-AI layouts (Carroll et al., 2019) and 32 layouts created via geometric transformations from a simple square base layout and randomly generated layouts to assess generalisation. The five original layouts enable comparisons to earlier work, and the transformed layouts expose whether the agent’s behaviour is robust (or ideally invariant) to layout transformations that do not require a different strategy. We secondly propose to evaluate in an ad-hoc teamwork setting (Stone et al., 2010), we train populations for 24 agents for the original five layouts via: Fictitious Co-Play (FCP) (Strouse et al., 2021) and Maximum Entropy Population Training (MEP) (Zhao et al., 2023). Each population includes low-, mid-, and high-skill checkpoints (10%, 50%, and 100% of achieved return) for diversity. MEP uses an entropy coefficient $\alpha=0.01$. This setting evaluates the adaptability of trained agents to diverse, possibly brittle partners. Finally, we also evaluate in a zero-shot coordination setting (Hu et al., 2020) in which we test an agent’s capability to adapt to partners that themselves were trained for the zero-shot coordination setting. Both ad-hoc teamwork and zero-shot coordination are evaluated on the original five.
We report two metrics. First, the mean episode reward and second, the solved rate: the proportion of episodes where at least two soups are delivered, distinguishing goal-directed behaviour from random actions. Finally, we also conduct a qualitative error analysis to examine failure modes across different levels.
Our primary evaluation metric based on episode return serves as a direct measure of both task efficiency and partner adaptability. Since rewards in Overcooked are only obtained by successfully delivering soups within a fixed time horizon, the return reflects how efficiently agents coordinate to complete tasks under temporal constraints. Additionally, measuring return in cross-play and ad-hoc teamwork settings isolates the agent’s ability to adapt to novel partners without prior joint training, thus capturing adaptability in zero-shot cooperation.
5 Experiments
@@
Figure 5: The SoftMoE-LSTM agents architecture used in this work. We employed the PerConv tokenisation technique introduced in Obando-Ceron et al. (2024).
Figure 6: Sample levels generated by the different methods after $15,000$ (Middle) and $30,000$ (End) epochs. Even after considerable training, none of the methods can guarantee the generation of solvable layouts (Middle row, leftmost and rightmost).
We conducted a series of experiments to establish performance baselines for partner and environment generalisation using state-of-the-art DCD methods and policy architectures. We first compare several policy architectures – CNN-LSTM, SoftMoE-LSTM, and S5-based models – on a fixed set of evaluation layouts to identify a strong baseline. We then assess whether agents trained with different DCD methods can generalise to novel environments without prior exposure to the test distribution. This is followed by experiments to establish coordination capabilities with unseen partners by testing agents in ad-hoc team play against diverse policy populations (FCP, MEP), as well as in zero-shot coordination with agents trained under the same method but different random seeds. Finally, we analyse failure modes through a detailed error analysis, identifying structural layout characteristics that correlate with poor performance, such as object spacing and path complexity.
All agents were trained using MAPPO (Yu et al., 2022), a strong cooperative multi-agent baseline. For DCD, we tested diverse algorithms with distinct generation principles (Dennis et al., 2020; Jiang et al., 2023): domain randomisation (DR), priority-based replay (robust parallel PLR⟂,∥), edit-based curriculum (parallel ACCEL), and learned environment design (Pop. PAIRED). We excluded POET (Wang et al., 2019) in this analysis as it outputs specialists rather than generalists, which we require (Parker-Holder et al., 2022). Additionally, we excluded MAESTRO as it is based on prioritised fictitious self-play (Heinrich et al., 2015; Vinyals et al., 2019) that is not readily adaptable to the cooperative setting (Strouse et al., 2021). We chose these methods as they offer better theoretical guarantees (PLR vs PLR), better runtime performance (ACCEL and PLR), or because we found them to perform better empirically (Pop. PAIRED vs PAIRED). To ensure a fair comparison, we standardised the environment design space: each method placed between one and 15 walls (either randomly or through a learned teacher policy), along with one or two items per object category (pots, onions, serving points, etc.). Layouts were procedurally generated using either teacher actions (PAIRED) or stochastic editing (ACCEL, PLR), with training conducted across five seeds, 32 parallel environments for 30,000 training iterations ( $\sim$ 400M steps). All architectural and training hyperparameters were selected via grid search and are detailed in Appendix A.4.
5.1 Experiment 1: Policies and Baselines
@@
Before evaluating generalisation on OGC, we identified a strong agent architecture that can serve as a policy backbone across all experiments. Comparing network architectures allows us to: (1) understand how architecture affects generalisation, and (2) establish upper-bound oracle performance on evaluation layouts, which will serve as reference points throughout the paper. We explored the following three architectures (see Appendix A.5 for details): CNN-LSTM is a standard convolutional encoder followed by an LSTM that demonstrated strong performance in previous work (Yu et al., 2023). SoftMoE-LSTM is an enhanced architecture using a Soft Mixture-of-Experts (SoftMoE) module (Obando-Ceron et al., 2024) and PerConv tokenisation, replacing the final layer of the CNN-LSTM. We explore SoftMoE agents because of their strong parameter scaling properties (Obando-Ceron et al., 2024). CNN-S5 is a CNN encoder paired with S5 layers (Smith et al., 2023) instead of LSTMs, inspired by structured state-space models (Gu et al., 2022) that showed strong performance in meta-RL (Lu et al., 2023). Each agent was trained on the five human-designed evaluation layouts (Cramped Room, Asymmetric Advantages, etc.) to assess whether the architecture could fit the joint task. In all experiments that follow, we refer to these as oracles – they have access to test environments during training and thus represent an empirical upper bound without generalisation.
As shown in Figure 7 (left), all architectures are capable of fitting the evaluation layouts in self-play. SoftMoE-LSTM achieves the highest returns across the board (Table 1), with lower variance and better stability. CNN-S5 significantly underperforms, suggesting S5 layers may not suit cooperative RL in this setting. To confirm these models did not simply memorise layout-specific strategies, we trained a SoftMoE-LSTM agent on a single layout and evaluated it on all five. The steep performance drop suggests overfitting, underscoring the importance of multi-layout training.
Key takeaway: We find that SoftMoE-LSTM generalises best among tested architectures, and adopt it for all subsequent experiments. This result also suggests that mixture-of-expert routing may support generalisation in multi-object, sparse-reward environments like Overcooked.
5.2 Experiment 2: Generalisation to Novel Environments
Table 1: Mean episode reward for the different methods averaged over the respective testing layouts. The best result is shown in bold. We report aggregate statistics over five random seeds and the 95% confidence interval. We include oracles trained on the five testing layouts to establish an empirical upper bound.
@@ 5.2 Experiment 2: Generalisation to Novel Environments
DR $2.84\pm 0.07$.37) $4.21\pm 0.09$.52) $0.06\pm 0.00$.17)
PLR $0.18\pm 0.00$.05) $0.68\pm 0.00$.75) $0.13\pm 0.00$.14)
PAIRED $0.52\pm 0.00$.39) $13.12\pm 0.11$.46) $0.02\pm 0.00$.44)
ACCEL $0.24\pm 0.00$.41) $0.61\pm 0.00$.55) $0.06\pm 0.00$.12)
Oracle $204.44\pm 0.29$.61) $216.76\pm 0.34$.41) $166.63\pm 0.23$.62)
@@
Table 2: Performance on mirrored, rotated and squeezed levels as illustrated in Figure 4. The transformations are grouped such that for each category, the optimal strategy remains the same modulo mirroring and rotation. Large, medium and small are defined in terms of the available movement space. Squeezed and squeezed small define increasingly narrow spaces. The large confidence interval suggests that agents do not learn general strategies for Overcooked. We analyse SoftMoE-LSTM agents.
Method Large Medium Small Squeezed Squeezed Small Average
DR $18.1\pm 3.0$.1) $13.1\pm 2.0$.2) $10.4\pm 1.5$.8) $11.3\pm 1.9$.9) $10.3\pm 1.7$.4) $13.1\pm 2.1$.1)
PLR $0.2\pm 0.0$.4) $0.7\pm 0.1$.0) $1.0\pm 0.0$.6) $0.0\pm 0.0$.0) $0.1\pm 0.0$.1) $0.5\pm 0.0$.5)
PAIRED $16.2\pm 1.7$.6) $31.5\pm 3.8$.4) $25.4\pm 2.8$.7) $4.7\pm 0.6$.9) $8.4\pm 0.8$.9) $19.9\pm 2.2$.2)
ACCEL $1.5\pm 0.2$.6) $1.4\pm 0.0$.6) $0.8\pm 0.0$.6) $0.2\pm 0.0$.4) $0.1\pm 0.0$.3) $0.9\pm 0.0$.8)
We then tested whether agents trained with unsupervised environment design can generalise to unseen environments. Unlike recent work that trained on augmented variants of test levels (Jha et al., 2025), agents in the OGC are faced with entirely new test environments without prior exposure.
As can be seen in Table 1, despite using tuned implementations and scalable architectures, most methods fail to generalise, achieving near-zero returns on the evaluation layouts. Only Population PAIRED has limited success, significantly outperforming other methods ($p<0.05$), with a mean solved rate of $14.6\%\pm 7.7$. All other methods barely go above $0\%$ solved levels, suggesting that training on randomly generated or edited layouts is insufficient to prepare agents for the coordination structure of evaluation tasks. We check and find that roughly a quarter of randomly generated training layouts are unsolvable, and many others require long, inefficient action sequences. This further highlights why these methods struggle: not only is the evaluation set challenging, but producing feasibly solvable training levels is itself a key difficulty. The SoftMoE-LSTM-PAIRED policy only shows mediocre performance on Asymmetric Advantages and Cramped Room and completely fails to coordinate effectively in more complex layouts, such as Counter Circuit or Forced Coordination. The training curve in Figure 8 confirms this: While the SoftMoE agent converges on training layouts, its generalisation gap on evaluation levels remains substantial and persistent.
To probe this further, we evaluate agents on systematically transformed versions of base layouts – including mirrored, rotated, and squeezed variants – that preserve the underlying coordination task but alter spatial structure. Ideally, a general policy should perform consistently across such transformations. However, performance fluctuates widely (see Table 2), revealing that agents fail to learn spatially invariant coordination strategies. This suggests that current methods often overfit to superficial spatial patterns rather than acquiring abstract cooperation skills. However, even PAIRED exhibits high variance across layouts, highlighting its limited robustness.
Key takeaway: Even state-of-the-art DCD methods fail to generalise to unseen, complex multi-agent coordination tasks. This suggests that the OGC introduces a richer and more challenging design space than prior UED environments and reveals limitations in current level generation strategies – highlighting the need for stronger curriculum learning and generalisation-aware training algorithms.
5.3 Experiment 3: Generalisation to Unknown Partners
@@
Figure 9: Ad-hoc team-play performance of SoftMoE-LSTM-PAIRED and other baselines with both an MEP and an FCP expert population. We measure the returns of multiple seeds across the five original layouts. Thinner lines on the bars indicate 95% confidence intervals around the means. While SoftMoE-LSTM-PAIRED outperforms simple baselines, it fails to reach oracle performance.
@@
Figure 10: Zero-shot cooperation returns of the oracle and the SoftMoE-LSTM-PAIRED agents (trained with different seeds). Each square shows the performance of the row and column seeds. Self-play performance is thus displayed on the diagonal.
To evaluate whether agents trained with DCD methods can coordinate with unknown partners in previously unseen environments, we investigated two settings: Ad-hoc teamplay and zero-shot coordination. We evaluated agents against diverse populations of pre-trained partners using two protocols: FCP (Strouse et al., 2021) and MEP (Zhao et al., 2023). Each population consisted of 24 agents with diverse skill levels and learning histories (see Appendix A.6.4).
Figure 9 shows the performance of the SoftMoE-LSTM-PAIRED agent compared to three baselines: a stationary partner (stay), a randomly acting agent (random), and an oracle trained on the evaluation layouts (see above). As can be seen from the figure, the SoftMoE agent consistently outperforms the baselines but fails to reach oracle performance. Notably, it often performs only slightly better than random coordination – a sign of poor robustness to novel partners. We hypothesise that this gap stems from the divergence between the training layouts (often open and simplified) and the evaluation layouts that have different cooperation demands. As shown in Figure 6, current DCD methods tend to converge toward minimal-complexity levels that facilitate early success but fail to expose agents to realistic partner dependencies.
We also assessed ZSC by evaluating whether the agent can coordinate with an independently trained copy of itself with a different random seed. This setting removes population diversity and isolates the agent’s generalisation to novel weights and latent partner strategies. As shown in Figure 10, SoftMoE-LSTM-PAIRED performs poorly across the more complex evaluation layouts (Coordination Ring, Forced Coordination, Counter Circuit), and even underperforms its oracle counterpart. Interestingly, in the simpler layouts, the agent occasionally achieves higher returns in cross-play than in self-play – suggesting that some diversity across seeds may help mitigate overfitting.
Key takeaway: DCD-trained agents – despite some progress in environment generalisation – struggle to generalise to novel partners in unknown environments.
5.4 Experiment 4: Qualitative Failure Analysis